OpenStudy (anonymous):

Since January 1, 1960, the population of Slim Chance has been described by the formula P=33000(0.98)^t where P is the population of the city t years after the start of 1960. At what rate was the population changing on January 1, 1978?

OpenStudy (anonymous):

Rate - so you want to find dP/dt when t=18 (1978-1960) \[\frac{dP}{dt}=ln(0.98)\cdot 33000(0.98)^t\] substitute t=18 to find your solution

OpenStudy (anonymous):

thank you!